OpenStudy (anonymous):

Help please!!!!!! What number Am I?! Clue 1: when you divide me by the number is 4 - Clue 2: I have 2digits both are odd. Clue 3- the sum of my digits is 10

jimthompson5910 (jim_thompson5910):

" when you divide me by the number is 4 " sounds clunky. Is there a typo here?

OpenStudy (cj49):

"when you divide me by the number is 4 " divide by wat ?

OpenStudy (anonymous):

no my cousin said that is what her teacher wrote on the board @cj49 @jim_thompson5910

OpenStudy (kittykitty04):

I think it's not divide... It might be subtract the numbers to get 4.

OpenStudy (kittykitty04):

When you look at clue 2 and clue 3. the sum is 10, you have either 3+7 or 1+9.

OpenStudy (kittykitty04):

so 7-3=4

OpenStudy (cj49):

the odd no are 1,3,5,7,9 and 7+3=10 9+1=10 if the first clue is that u need to subtact the number then either the no is 73 or 37

OpenStudy (kittykitty04):

Otherwise, I think I'm stuck.

OpenStudy (anonymous):

i got 7+3 as well but the divide 4 threw me off

OpenStudy (anonymous):

@kittykitty04

OpenStudy (anonymous):

@cj49

OpenStudy (anonymous):

is it 73 or 37 @ kittykitty04 @cj49

OpenStudy (cj49):

it should be the difference between the two no is 4...and it can either be 37 or 73